OpenStudy (anonymous):

Prove (sec x - tan x)^2 = (1-sin x)/1+sin x

OpenStudy (anonymous):

I distribute the square of 2 first right?

OpenStudy (solomonzelman):

I would think so... yes.

OpenStudy (anonymous):

1 - sin ^2 x = cos ^2 x?

OpenStudy (solomonzelman):

what is this for here, right now?

OpenStudy (anonymous):

what you mean?

OpenStudy (anonymous):

oh

OpenStudy (solomonzelman):

no, they are not squared, my bad...

OpenStudy (anonymous):

do you still need help

OpenStudy (solomonzelman):

yes, just distribute the power of 2....

OpenStudy (anonymous):

then i get \[\frac{ 1 }{ \cos^2 x } - \frac{ \sin^2 x}{ \cos^2 x}\]

OpenStudy (anonymous):

thats after simplfying from sec^2 x - tan ^2 x

OpenStudy (solomonzelman):

I don't see a need in doing it NOW.